# Static or Dynamic: Towards Query-Adaptive Token Selection for Video Question Answering

### v1

@@ Appendix A Prompt Details
We utilize the template provided by the model for the instruction prompt part. We only introduce the textual organization format in the questioning part.
A.1 Prompts for Multiple-Choice Questions
@@
And here is an example for EgoSchema:
@@
Identify the recurring actions in the video and briefly discuss their significance to the overall narrative.
A. C constantly organizing a plastic box, suggesting her obsession with tidiness
B. C and the boy taking turns throwing objects out of the window, showcasing a game
C. C pouring water and conversing with the boy, highlighting routine and communication
D. The boy trying to get C’s attention by throwing a toy on the blanket repeatedly
E. C teaching the boy how to fold blankets properly and arrange his toys
A.2 Prompts for Open-Ended Questions
We add the sentence "Answer the question according to the video." at the beginning of the multiple-choice questions. Here is an example for questions in MSVD-QA:

### v2

@@ Appendix B Prompt Details
We use the template provided by the model for the instruction prompt, while introducing our textual organization format only in the questioning part.
B.1 Prompts for Multiple-Choice Questions
@@
Here is an example for EgoSchema:
@@
Taking into account all the actions performed by c, what can you deduce about the primary objective and focus within the video content?
A. C is cooking.
B. C is doing laundry.
C. C is cleaning the kitchen.
D. C is cleaning dishes.
E. C is cleaning the bathroom.
And here is an example for MLVU:
Respond with only the letter (A, B, C, D, E or F) of the correct option.
In what setting does the video take place?
(A) Castle
(B) Forest
(C) Desert
(D) Countryside
(E) Ocean
(F) Campus
B.2 Prompts for Open-Ended Questions
We add the sentence "Answer the question according to the video." at the beginning of the open-ended questions. Here is an example: